A certain body weighs 12N in air and 10 N in when immersed in water. Find its

total volume and its density.


1
Expert's answer
2022-03-03T12:04:51-0500

The weight of body in water

"W=mg-g\rho_wV"

Hence, the volume of a body

"V=\frac{mg-W}{g\rho_w}=\frac{12-10}{9.8*1000}=2*10^{-4}\:\rm m^3"

The density of a body

"\rho=\frac{m}{V}=\frac{12}{9.8*2*10^{-4}}=6000\:\rm kg /m^3"